OpenStudy (anonymous):

A cyclist rides her bike at a rate of 9 kilometers per hour. What is this rate in kilometers per minute? How many kilometers will the cyclist travel in 2 minutes? Do not round your answers.

OpenStudy (anonymous):

First, close your mouth. Second, how many minutes are in an hour?

OpenStudy (anonymous):

60 minutes.

OpenStudy (anonymous):

Good, now how would you express 9 kilometers per hour in short form?

OpenStudy (anonymous):

9 km?

OpenStudy (anonymous):

You forgot the hour Anyway it would be 9km/hour So something like this: \[9\frac{ km }{ hr }\] You have to change it into minutes so you have to eliminate the hour How would you do that? Take a guess

OpenStudy (anonymous):

Would you multiply it by itself?

OpenStudy (anonymous):

Good! That is correct. You will multiply it. So how would we write the fraction we will multiply such that: \[9\frac{ km }{ hr } \times \frac{ hr }{\min} = 9\frac{ km }{ \min }\]

OpenStudy (anonymous):

Im sorry, disregard the 9 after the equal sign

OpenStudy (anonymous):

This is the correct one. \[9\frac{ km }{ hr }×\frac{ hr }{ \min }= ?\frac{ km }{ \min }\]

OpenStudy (anonymous):

Is that all needed for the rate? Would we have to divide 9 by 60 mins?

OpenStudy (anonymous):

Yep so you multiply it by 1/60 to find the rate

OpenStudy (anonymous):

3/20? and in decimal form its .15?

OpenStudy (anonymous):

That's right Now find out how many km did the cyclist travel in two minutes

OpenStudy (anonymous):

multiplying the rate by 2 isn't going to give me the answer right?

OpenStudy (anonymous):

nope, i think you gotta divide it

OpenStudy (anonymous):

But it doesnt make sense for some reason

OpenStudy (anonymous):

I divided 3/20 and got 1/60

OpenStudy (anonymous):

I think you should multiply it by 2

OpenStudy (anonymous):

.3?

OpenStudy (anonymous):

That is what i am thinking, but i need to confirm. Give me a second

OpenStudy (anonymous):

0.3 is correct. You can do it this way 2min of 1 hour is 1/30 hr divide 9/30 You get 0.3km

OpenStudy (anonymous):

Thank you for confirming it and helping me out!!

OpenStudy (anonymous):

No problem!